# Pre-Processing Matters: Similarity-based Simulated Annealing Approach for Fast Parameter Tuning in WSI Classification

### v2

@@ Introduction
For single-domain inference performance, our experiments demonstrate that pre-processing parameters yielding accurate predictions produce images with similar structures (as illustrated in Figure 5). Therefore, we only evaluate the pre-processed Whole Slide Images (WSIs) that closely resemble the optimally pre-processed WSI. To quantitatively assess differences between two sets of pre-processing results, we use Peak Signal-to-Noise Ratio (PSNR) to measure the similarity of images generated during patch creation. As depicted in Figure 1, our workflow includes creating patches, extracting features, training models, and testing inference data. During patch generation, thumbnails are created for visualizing the preprocessing effects. Thus, we can determine the similarity between these two sets of pre-processed results by calculating PSNR for these thumbnails.
PSNR
@@
[…]
In our motivation experiments, which explore the impact of pre-processing parameters on out-of-domain performance, we have already processed sufficient Whole Slide Images (WSIs) with inference results. We established the similarity threshold $\tau$ based on previous experimental analyses. For our proposed SSAPT method, we use ${H_{train}}$ as the default parameter. Using the current optimal parameter $H^{*}$, we generate a new parameter ${H_{new}}$, create patches, and calculate their PSNR against the optimal thumbnails $I^{*}$. If the PSNR exceeds threshold $\tau$, we proceed to more resource-intensive tasks such as feature extraction in histopathological analysis and subsequent inference using an existing model. Each new parameter set ${H_{new}}$ is added to our historical dataset $R=\{(H_{1},y_{1}),...,(H_{i},y_{i})\}$. If ${H_{new}}$ demonstrates improved performance, there is a probability $p$ (as per Eq.(3)) that ${H_{*}}$ will be updated with ${H_{new}}$. This process will be further discussed in the following subsection.
Similarity-based Simulated Annealing for Parameter Tuning
Algorithm 1 Similarity based Simulated Annealing for Parameter Tuning (SSAPT)
Input: Default parameter ${H_{train}}$ with result set $\mathcal{R}$, Hyper-parameters space $\Phi$, initial temperature $T_{0}$, minimum temperature $T_{\text{min}}$, cooling rate $\alpha$, similarity threshold $\tau$, function $Infer$ to evaluate performance, function $S$ to calculate similarity, $p$ is the acceptance probability function
@@
2:  Generate pre-processed thumbnail $I^{*}$ from $H^{*}$
@@
6:  $T\leftarrow T_{0}$
7:  while $T>T_{\text{min}}$ do
8:     Select $H_{\text{new}}$ by perturbing $H^{*}$ slightly from $\Phi$
9:     if  $H_{\text{new}}$ not in $\mathcal{R}$ then
10:        Generate thumbnail $I_{\text{new}}$ with $H_{\text{new}}$
11:        $\Delta s\leftarrow S(I_{\text{new}},I^{*})-\tau$
12:        if $\Delta s>0$  then
13:           Extract feature $F_{\text{new}}$      $\triangleright$ Expensive step
14:           $y_{\text{new}}\leftarrow Infer(F_{\text{new}})$
15:           $\mathcal{R}\leftarrow\mathcal{R}\cup(H_{\text{new}},y_{\text{new}})$
16:           if $y_{\text{new}}>y^{*}$  then
17:              $H^{*},I^{*},y^{*}\leftarrow{H_{\text{new}}},I_{\text{new}},y_{\text{new}}\text{ with $p$}$
@@
19:        end if
20:        $T\leftarrow\alpha T$
21:     end if
22:  end while
23:  return  $H^{*}$
To avoid falling into local optima, we integrate the Simulated Annealing algorithm (Kirkpatrick, Gelatt Jr, and Vecchi 1983) into our method to identify optimal parameters. We detail SSAPT in Algorithm 1. Our experimental results indicate that creating patches with thumbnails and bypassing most parameter feature extraction processes allows us to refine pre-processing parameters for enhanced inferential performance.
Our algorithm is designed to optimize pre-processing parameters effectively, ensuring high performance while accelerating the overall process. The Simulated Annealing algorithm prevents convergence at local optima by exploring multiple random states near the current solution, akin to simulated annealing itself. Occasionally, we enhance solutions by leveraging the current temperature to explore nearby states post-annealing. This approach introduces randomness through annealing rules, significantly increasing the likelihood of achieving an optimal solution tailored for histopathological tasks.
During each iteration, we aim to match the pre-processed results closely with the current optimum. We use a similarity threshold $\tau$, related to PSNR as a filter mechanism. After patch creation, features are extracted and analyzed only from pre-processed Whole Slide Images (WSIs) that exhibit a PSNR higher than $\tau$. WSIs with lower PSNR compared to the optimal pre-processed WSI are excluded from further processing. If superior inference performance emerges, updates are made for $H^{*},I^{*},y^{*}$ based on acceptance probability $p$:
$p(\Delta s)=\exp(\Delta s/T)$ (3)
Experiments
Dataset Center ABMIL DSMIL CLAM Bayes-MIL MHIM-DSMIL
Default Ours Default Ours Default Ours Default Ours Default Ours
Camelyon17 C0 0.93 0.97 0.77 0.96 0.921 0.95 0.83 0.94 0.87 0.96
C1 0.87 0.92 0.85 0.89 0.512 0.846 0.35 0.8 0.81 0.87
C4 0.88 0.93 0.92 0.96 0.874 0.92 0.85 0.92 0.91 0.94
TCGA-COAD CM 0.75 0.778 0.75 0.778 0.75 0.889 0.861 0.889 0.75 0.801
D5 0.709 0.774 0.806 0.871 0.742 0.871 0.709 0.806 0.839 0.871
DM 0.783 0.826 0.869 0.869 0.739 0.739 0.783 0.826 0.826 0.87
G4 0.963 0.963 0.963 0.963 0.889 0.926 0.926 1 0.925 0.963
TCGA-BRCA A7 0.902 0.961 0.921 0.941 0.941 0.961 0.882 0.882 0.923 0.961
AC 0.775 0.8 0.725 0.8 0.75 0.8 0.775 0.775 0.75 0.775
AR 0.781 0.8 0.765 0.781 0.813 0.828 0.813 0.828 0.75 0.766
C8 0.907 0.93 0.581 0.813 0.907 0.93 0.86 0.884 0.907 0.93
@@
Table 2: Experiments in out-of-domain inference performance. We compare the accuracy obtained by default hyper-parameters with the optimal hyper-parameter searched by our SSAPT algorithm.
Experiments Setup
Our experiment utilizes the pre-processing pipeline from CLAM. The datasets employed in this study include Camelyon 16, Camelyon 17 (Litjens et al. 2018), TCGA-NSCLC, TCGA-BRCA, and TCGA-COAD (Weinstein et al. 2013). For Camelyon16 and Camelyon17, we focus on normal-tumor classification; for all TCGA datasets, we address subtype tasks.
@@
In our experiments, we standardize seeds and model hyperparameters such as learning rate and loss weight across similar tasks within the same model framework. Additional experimental details are provided in the Appendix. We employ K-fold, StratifiedKFold or k-fold Monte Carlo cross-validation methods (with k set to 10) to train models and assess out-of-domain performance. All results presented in Table 2 and Table 3 represent average accuracies calculated over these 10-fold models.
Classification Accuracy Validation of SSAPT
@@
For the Camelyon datasets, we trained our model using 270 whole slide images (WSIs) from Camelyon16 and tested inference performance across various centers in Camelyon17 to simulate out-of-domain scenarios. It’s important to note that Camelyon16 data was collected from UMCU and RUMC, while Camelyon17 includes data from CWZ, RST, UMCU, RUMC, and LPON. Consequently, only centers 0 (CWZ), 1 (RST), and 4 (LPON) of Camelyon17 represent out-of-domain situations.
@@
We compared the inference performance using features extracted by default parameters (not the worst) and those optimized through our SSAPT algorithm. As shown in Table 2, the optimal hyper-parameters identified by our SSAPT algorithm generally surpass the default settings across various models and datasets. In specific cases, such as with CLAM and BayesMIL at Center 1, performance improved dramatically from below 0.5 to up to 0.8 simply by adjusting the hyper-parameters. Even for robust models like ABMIL, modifying parameters proved effective.
For TCGA datasets, we divide the entire dataset according to TCGA Tissue Source Site Codes. We select several centers with a sufficient number of WSIs to form the testing sub-dataset, while the remaining data constitutes the training dataset. The out-of-domain inference performance also improves (as shown in Table 2), indicating that our issue is general and our SSAPT provides a universal solution. Additionally, we conduct experiments on a larger TCGA-NSCLC dataset that includes more extensive out-of-domain testing (as detailed in Table 3).
Dataset Center ABMIL CLAM Bayes-MIL
Def. Ours Def. Ours Def. Ours
TCGA- NSCLC 21 0.722 0.778 0.556 0.778 0.556 0.667
22 0.865 0.892 0.784 0.838 0.838 0.838
43 0.750 0.800 0.600 0.750 0.700 0.800
49 0.595 0.643 0.690 0.714 0.714 0.762
55 0.495 0.516 0.659 0.758 0.275 0.330
77 0.932 0.932 0.909 0.932 0.750 0.800
97 1.000 1.000 0.913 1.000 0.869 0.957
Table 3: Comparison of inference accuracy in TCGA-NSCLC. We evaluate the default hyper-parameter with optimal hyper-parameter searched by our SSAPT Algorithm in a larger dataset TCGA-NSCLC with more centers.
Efficiency Validation
Figure 4 shows the running time for tuning preprocessing parameters 100 times. Our Simulated Annealing algorithm for Parameter Tuning uses PSNR to compare each pre-processed WSI with the best one so far, selecting only those with high similarity for further feature extraction. WSIs with low PSNR are skipped, streamlining the process and ensuring that only promising candidates are evaluated further. This approach prevents our SSAPT from getting stuck in local optima and saves time by reducing unnecessary feature extractions.
Ablation Study
PSNR
Without the PSNR-based strategy, the algorithm performs a random search. It randomly generates pre-processing parameters, creates patches, extracts features, and assesses performance in each iteration without skipping feature extraction (as shown in Table 4).
Figure 5: The evidence of PSNR-based mechanism works in our fast hyper-parameter search in histopathological tasks.
We also assess FDD-Simulated Annealing. It’s important to note that the Fréchet domain distance (FDD) in this part of the ablation study serves a different role compared to its use in the Discussion section. Here, FDD replaces PSNR for comparison purposes. Both FDD and PSNR-based mechanisms in this ablation study are employed to constrain the search space for optimal hyper-parameters, rather than searching across the entire hyper-parameter space.
Table 4: Importance of the PSNR. Results are based on optimizing 100 times in the center1 of Camelyon17. SA stands for Simulated Annealing, Lat. stands for Latency and Mem. stands for Memory.
Strategy Acc Lat. (min) Mem. (GB)
None (Baseline) 0.847 9600 1250
FDD-SA 0.846 5350 650
PSNR-SA (Ours) 0.846 1770 170
Simulated Annealing
@@
Simulated Annealing prevents optimization from getting stuck in local optima. As illustrated in Table 5, search algorithms lacking Simulated Annealing repeatedly test hyper-parameters that yield a higher PSNR than the current optimal solution.
Table 5: Importance of Simulated Annealing (SA). Results based on optimizing 100 times in center1 of Camelyon17.
Strategy Accuracy Latency (min)
Without SA 0.829 4119
With SA 0.846 1770
Discussion
Quantifying the Initial Features
We demonstrate that significant differences exist in the initial features extracted by various pre-processing parameters, which are used as inputs for training models or making inferences. These differences arise not only from distribution shifts but also from variations in the number of patches extracted from each WSI, leading to partial information loss. For the same WSI, we compare features generated under different hyper-parameter settings: $H_{opt}$, $H_{good}$, and $H_{random}$. As shown in Figures 1 and  6, features are similar when their inference performance is comparable. However, Figure 6 (Right) highlights distinct differences between $H_{opt}$ and $H_{random}$ within the t-SNE space, underscoring how patch count variations can exacerbate feature representation discrepancies and affect out-of-distribution (OOD) inference performance.
Figure 6: The visualization of compasion between features. Left: The comparison of the feature obtaining high accuracy with the optimal feature. Right: The comparison of the feature extracted by random hyper-parameters with the optimal one. Note that these features are extracted from the same WSIs.
Evaluating the Influence of Pre-processing on Model Representations
We utilize Fréchet domain distance (FDD) (Pocevičiūtė et al. 2023) to compare features from the penultimate layer and assess how pre-processing parameters influence the model’s representations. For each specific out-of-domain testing dataset $D_{val}^{i}$, WSIs undergo pre-processing using $H_{i}$ to generate feature $F_{i}$. This feature is then input into the model for inference. We extract the penultimate layer feature $P_{i}$ before it reaches the final classifier. The same process applies when using optimal pre-processing features identified by our SSAPT algorithm, denoted as $H_{opt}$, resulting in $F_{opt}$. For WSIs processed with $H_{i}$, a higher accuracy correlates with a lower FDD compared to those processed with optimal preprocessing ($H_{opt}$), as illustrated in Figure 7.
Figure 7: Quantifying the Features Impact by Model Visibility.
SSAPT is designed for histopathological tasks
Previous hyper-parameter optimization stategies (Falkner, Klein, and Hutter 2018; Choi et al. 2024) have been developed with a focus on general machine learning models and evaluated on datasets characterized by a large number of images but relatively small image sizes. However, the unique characteristics of WSIs should not be ignored. WSIs are giga-pixel images that demand significantly higher computational resources. Despite advancements in hardware, memory constraints still pose a significant challenge in the digital processing of WSIs in current research (Cifci et al. 2023; Echle et al. 2021). Additionally, nature images have very different data distribution compared to the whole slide image (Raghu et al. 2019). Therefore, designing hyper-parameter optimization stategy for CPath is necessary. The proposed SSAPT algorithm is the first hyper-parameter optimization stategy designed for histopatholocial tasks.
Conclusion

### v3

@@ 4.3 Optimization of Fast Hyper-Parameter Search
Therefore, we only evaluate the pre-processed Whole Slide Images (WSIs) that closely resemble the current optimally pre-processed WSI in each optimization iteration. To quantitatively assess differences between two sets of pre-processing results, we use Peak Signal-to-Noise Ratio (PSNR) to measure the similarity of images generated during patch creation. The general histopathological workflow includes creating patches, extracting features, training models, and testing inference data. During patch generation, thumbnails are created for visualizing the preprocessing effects. Thumbnails are 64x downsampled from original giga-pixel WSIs. Thus, we can determine the similarity between these two sets of pre-processed results by calculating PSNR for these thumbnails efficiently.
4.3.2 Peak Signal-to-Noise Ratio (PSNR)
@@
4.3.3 PSNR-based Basin Hopping for Hyper-Parameter Optimization
Algorithm 1 PSNR-based Basin Hopping Hyper-Parameter Optimization (BAHOP)
Input: Default parameter ${H_{train}}$ with result set $\mathcal{R}$, Hyper-parameters space $\Phi$, initial similarity threshold $\tau$, function $Infer$ to evaluate performance, function $S$ to calculate similarity
@@ 4.3.3 PSNR-based Basin Hopping for Hyper-Parameter Optimization
2:  Generate pre-processed thumbnails $I^{*}$ from $H^{*}$ (for all validation WSIs)
@@
6:  while optimization times do
7:     Select $H_{\text{new}}$ by perturbing $H^{*}$ slightly from $\Phi$
8:     if  $H_{\text{new}}$ not in $\mathcal{R}$ then
9:        Generate thumbnails $I_{\text{new}}$ with $H_{\text{new}}$
10:        if $S(I_{\text{new}},I^{*})>\tau$  then
11:           Extract feature $F_{\text{new}}$      $\triangleright$ Expensive step
12:           $y_{\text{new}}\leftarrow Infer(F_{\text{new}})$
13:           $\mathcal{R}\leftarrow\mathcal{R}\cup(H_{\text{new}},y_{\text{new}})$
14:           if $y_{\text{new}}>y^{*}$  then
15:              $H^{*},I^{*},y^{*}\leftarrow{H_{\text{new}}},I_{\text{new}},y_{\text{new}}$
16:           end if
17:        end if
@@
19:  end while
20:  return  $H^{*}$
Basin-hopping is a global optimization technique characterized by iterative processes involving random perturbation of coordinates, followed by localized optimization and subsequent evaluation and acceptance or rejection of new coordinates based on minimized function values. This method is particularly useful in high-dimensional landscapes. However, an inherent limitation of Basin-hopping lies in the necessity to execute random perturbations at a designated point, typically a local minimum. These perturbations must be suitably sized - adequately large to escape the current local minimum yet not excessively to devolve into total randomness. To address this challenge, we introduce a Peak Signal-to-Noise Ratio (PSNR) threshold within the Basin-hopping optimization process. Subsequently, our experimental endeavors focus on establishing a correlation between pre-processing methodologies applied to whole slide images and the PSNR metric.
The integration of PSNR-based filtering does not fundamentally change the Basin Hopping process itself but rather serves as a heuristic pre-processing step to enhance efficiency and fit histopathological tasks. The PSNR threshold is set dynamically and adapts across different datasets. The experiments of the effection of each pre-processing parameter are shown in the Appendix. From the experiments (as shown in the Appendix), the criteria for selecting the PSNR threshold and dynamically adapting to different datasets is that we perturb the pre-processing parameter – segmentation threshold (related to foreground and background detection). Each time, we change the value to plus or minus 1. We calculate the PSNR for this comparison as the initial PSNR. This process only requires the creation of a patch and quick PSNR calculation in thumbnails.
Also, we delete the mechanism of temperature and accepted probability from the original Basin hopping, ensuring our BAHOP is designed properly for histopathological tasks. For our proposed BAHOP method (as shown in Algorithm 1), we use ${H_{train}}$ as the default parameter. Using the current optimal parameter $H^{*}$, we generate a new parameter ${H_{new}}$, create patches, and calculate their PSNR against the optimal thumbnails $I^{*}$. If the PSNR exceeds threshold $\tau$, we proceed to more resource-intensive tasks such as feature extraction in histopathological analysis and subsequent inference using an existing model. Each new parameter set ${H_{new}}$ is added to our historical dataset $R=\{(H_{1},y_{1}),...,(H_{i},y_{i})\}$. If ${H_{new}}$ demonstrates improved performance, ${H_{*}}$ will be updated with ${H_{new}}$.
5 Experiments
Model TCGA-COAD TCGA-BRCA
Center Accuracy AUC Center Accuracy AUC
Default Ours Default Ours Default Ours Default Ours
ABMIL [12] CM 0.75 0.778 0.828 0.838 A7 0.902 0.961 0.964 0.994
D5 0.709 0.774 0.64 0.647 AC 0.775 0.805 0.728 0.761
DM 0.783 0.826 0.732 0.759 AR 0.781 0.816 0.834 0.856
G4 0.963 0.963 0.98 1 C8 0.907 0.93 0.867 0.908
DSMIL [15] CM 0.75 0.778 0.848 0.859 A7 0.921 0.941 0.923 0.962
D5 0.806 0.871 0.5 0.493 AC 0.725 0.805 0.761 0.761
DM 0.869 0.869 0.777 0.804 AR 0.765 0.781 0.733 0.738
G4 0.963 0.963 0.96 0.98 C8 0.581 0.813 0.767 0.75
CLAM [17] CM 0.75 0.889 0.56 0.636 A7 0.941 0.961 0.947 0.974
D5 0.742 0.871 0.46 0.793 AC 0.75 0.81 0.772 0.761
DM 0.739 0.739 0.571 0.571 AR 0.813 0.828 0.882 0.878
G4 0.889 0.926 0.94 0.96 C8 0.907 0.93 0.867 0.875
BayesMIL [5] CM 0.861 0.889 0.929 0.929 A7 0.882 0.882 0.934 0.955
D5 0.709 0.806 0.667 0.593 AC 0.775 0.775 0.799 0.821
DM 0.783 0.826 0.821 0.875 AR 0.813 0.828 0.796 0.791
G4 0.926 1 0.96 1 C8 0.86 0.884 0.792 0.767
MHIM- DSMIL  [23] CM 0.75 0.801 0.77 0.798 A7 0.923 0.961 0.957 0.996
D5 0.839 0.871 0.48 0.493 AC 0.75 0.775 0.611 0.745
DM 0.826 0.87 0.714 0.786 AR 0.75 0.766 0.709 0.713
G4 0.925 0.963 0.94 0.98 C8 0.907 0.93 0.742 0.733
@@ 5 Experiments
Table 2: Experiments about out-of-domain inference performance in TCGA-COAD and TCGA-BRCA for cancer subtype. We compare the accuracy obtained by default hyper-parameters with the optimal hyper-parameter searched by our BAHOP algorithm.
Model Out-of-Domain(Center 0) Out-of-Domain(Center 1) Out-of-Domain(Center 4)
Accuracy AUC Accuracy AUC Accuracy AUC
Def Ours Def Ours Def Ours Def Ours Def Ours Def Ours
ABMIL [12] 0.93 0.97 0.921 0.965 0.84 0.92 0.833 0.896 0.88 0.93 0.907 0.911
DSMIL [15] 0.77 0.96 0.88 0.937 0.64 0.89 0.721 0.93 0.92 0.96 0.952 0.947
CLAM [17] 0.92 0.95 0.919 0.963 0.512 0.846 0.708 0.833 0.874 0.92 0.904 0.941
Bayes-MIL [5] 0.83 0.94 0.956 0.884 0.35 0.80 0.819 0.881 0.85 0.92 0.931 0.899
MHIM-DSMIL [23] 0.87 0.96 0.911 0.956 0.77 0.87 0.85 0.921 0.91 0.94 0.949 0.955
Table 3: Out-of-domain inference performance in Center 0, center1 and Center 4 of Camelyon17 for normal and tumor classification.
5.1 Experiments Setup
@@
Our experiment utilizes the pre-processing pipeline from CLAM [17]. The datasets include Camelyon 16, Camelyon 17 [16], TCGA-NSCLC, TCGA-BRCA, and TCGA-COAD [29]. For Camelyon16 and Camelyon17, we focus on normal-tumor classification; for all TCGA datasets, we address cancer subtype tasks. In our experiments, we fix the seeds and model hyperparameters such as learning rate and loss weight across similar tasks within the same model framework. Additional experimental details are provided in the Appendix. We employ K-fold, Stratified K-Fold or K-fold Monte Carlo cross-validation methods (with k sets to 10) to train models and assess out-of-domain performance. All results presented in Table 2, Table 3 and Table 4 represent average accuracy calculated over 10-fold models.
5.2 Classification Accuracy Validation of BAHOP
@@
For the Camelyon datasets, we use 270 WSIs from Camelyon 16 to train the model and test inference performance across various centers in Camelyon17 to simulate out-of-domain scenarios. It’s important to note that Camelyon16 data was collected from two centers (UMCU and RUMC), while Camelyon17 includes data from five centers (CWZ, RST, UMCU, RUMC, and LPON). Consequently, only centers 0 (CWZ), center 1 (RST), and center 4 (LPON) of Camelyon17 represent out-of-domain situations.
We compare inference performance using features extracted by default parameters (not the worst) and those optimized through our BAHOP algorithm. As shown in Table 2 and Table 3, the optimal hyper-parameters identified by our BAHOP algorithm generally surpass the default settings across various models and datasets. In specific cases, such as CLAM and BayesMIL at Center 1 of Camelyon 17, performance improved dramatically from 0.512 to 0.847 simply by adjusting the hyper-parameters.
For TCGA datasets, we divide the entire dataset according to TCGA Tissue Source Site Codes [29]. We select several centers with sufficient WSIs to form the testing sub-dataset, while the remaining data constitutes the training dataset. The out-of-domain inference performance also improves (as shown in Table 2), indicating that our issue is general and our BAHOP provides a universal solution. Additionally, we conduct experiments on a larger TCGA-NSCLC dataset that includes more extensive out-of-domain testing (as detailed in Table 4).
TCGA-NSCLC
Center ABMIL [12] CLAM [17] Bayes-MIL [5]
Def. Ours Def. Ours Def. Ours
21 0.72 0.78 0.56 0.78 0.56 0.67
22 0.87 0.89 0.78 0.84 0.84 0.84
43 0.75 0.80 0.60 0.75 0.70 0.80
49 0.59 0.64 0.69 0.71 0.71 0.76
55 0.49 0.52 0.66 0.76 0.28 0.33
77 0.93 0.93 0.91 0.93 0.75 0.80
97 1.00 1.00 0.91 1.00 0.87 0.96
Table 4: Comparison of inference accuracy in TCGA-NSCLC with more centers. Only accuracy is reported since each separate center only has one category. Def. stands for default parameter.
5.3 Comparison of other hyper-parameter optimization
As illustrated in Table 5, we compare different hyper-parameter optimization (HPO) techniques in one center of Camelyon 17 over 100 iterations. Our BAHOP can find the optimal hyper-parameters much quicker compared to alternative HPO methodologies. While other HPO techniques demonstrated the capability to achieve commendable accuracy levels, they incurred substantial costs in terms of computational time and resource utilization.
Table 5: Comparison of other hyper-parameter optimization. SA stands for Simulated Annealing. Lat. stands for Latency (smaller is better) and Mem. stands for Memory.
Strategy Acc Lat.(min) Mem.(GB)
Random search 0.845 9600 1250
Grid search 0.847 9600 1250
SA [27] 0.846 9600 1250
Bayes OPT [22] 0.834 9600 1250
BAHOP (Ours) 0.846 1770 170
5.4 Efficiency Validation
Figure 6 shows the running time for optimizing preprocessing parameters 100 times. Our PSNR-based Basin Hopping (BAHOP) algorithm for Parameter Tuning uses PSNR to compare each pre-processed WSIs with the best one so far, selecting only those with high similarity for further feature extraction. WSIs with low PSNR are skipped, streamlining the process and ensuring that only promising candidates are evaluated further. This approach prevents our BAHOP from getting stuck in local optima and saves time by reducing unnecessary feature extractions.
@@ 5.4 Efficiency Validation
5.5 Ablation Study
Strategy 1: PSNR
The workflow of inference in histopathological tasks includes creating patches, extracting features, and inferred in a pre-trained model. Without the PSNR-based strategy, the algorithm will run the most expensive step – feature extraction, for each iteration. Our PSNR machanism can skip feature extraction in many iterations (as shown in Table 6).
Strategy 2: Basin Hopping
@@ 5.5 Ablation Study
Basin Hopping prevents optimization from getting stuck in local optima. As illustrated in Table 5, search algorithms lacking Basin Hopping repeatedly test hyper-parameters that yield a higher PSNR than the current optimal solution. Since other optimizaion techniques have optimized for 100 times, the optimal accuracy is all high but the costs are so huge both in time and disk space.
Table 6: Ablation Study of the PSNR and Basin Hopping. Results are based on optimizing 100 times in the center1 of Camelyon 17. Lat. stands for Latency (smaller is better) and Mem. stands for Memory.
PSNR BH Acc Lat. (min) Mem. (GB)
0.846 9600 1250
0.845 2553 278
0.846 1770 170
6 Conclusion